# Quantum State Preparation for Probability Distributions with Mirror Symmetry Using Matrix Product States

### v2

@@ 4.2 Future Work
We have proposed a method for generating higher-accuracy quantum circuits from matrix product states by utilizing the mirror symmetry of the normal distribution and other functions. In this work, we focused on partially monotonically increasing functions, but extending this approach to other types of functions is a task for the future. Additionally, the precision of our method is contingent upon the accuracy of the conversion from tensor networks to quantum circuits. Hence, proposing methods for more precise conversion is a significant challenge. As part of this challenge, it would be intriguing to investigate whether incorporating machine learning [11, 12, 18, 32, 33] into our current method could further enhance accuracy.

### v3

@@ IV.3 Future Work
We have proposed a method for generating higher-accuracy quantum circuits from matrix product states by utilizing the reflection symmetry of the normal distribution and other functions. In this work, we focused on partially monotonically increasing functions, but extending this approach to other types of functions is a task for the future. Additionally, the precision of our method is contingent upon the accuracy of the conversion from tensor networks to quantum circuits. Hence, proposing methods for more precise conversion is a significant challenge. As part of this challenge, it would be intriguing to investigate whether incorporating machine learning [11, 12, 18, 40, 41] into our current method could further enhance accuracy.